Q1. A gaseous mixture contains 10gram of H2. 196g of nitrogen 192 gram of
oxygen. Pressure of hydrogen is 1.9atm. calculate the partial pressure of oxygen

gas and nitrogen gas.
1
Expert's answer
2020-12-07T03:19:08-0500

"P=n(" "RT\over V" ")" "-n_{constant}"

Moles of "N_2=n_{n_2}" "=196g\over 28g/mol" "=7mol"

Moles of "O_2=n_{O_2}=" "192g\over 32g/mol" "=6mol"

From the ideal gas law;

"P_{N_2}=" "n_{N_2}RT\over V"

"=" "{7mol\times 0.08206{atm.L\over mol.K} \times 273K}\over1L" "=156.82atm"

"P_{O_2}=" "{6mol\times 0.08206{atm.L\over mol.K} \times 273K}\over 1L" "=134.0atm"

The partial pressure of "N_2" "=156.82atm"

The partial pressure of"O_2=134atm"